OpenStudy (anonymous):

Compare: 3/4 { } 9/12 < > = None of the above

OpenStudy (anonymous):

OK how do you simplify a fraction?

OpenStudy (zzr0ck3r):

\[\frac{3}{4}*\frac{3}{3}=\frac{?}{?}\]

OpenStudy (anonymous):

You divide both 9 and 12 by three which gives you 3 and 4.

OpenStudy (anonymous):

So 9/12 simplified is 3/4, they are equal.

OpenStudy (anonymous):

thanks i had the answer i just wanted to know if i was right =D

OpenStudy (anonymous):

your welcome :)

OpenStudy (zzr0ck3r):

\[\frac{3}{4}=\frac{3}{4}*1=\frac{3}{4}\frac{3}{3}=\frac{9}{12}\\note\\\frac{3}{3}=1\]

OpenStudy (anonymous):

zzr0ck3r how do you figure that?